### v1

# Improving and Evaluating Open Deep Research Agents

Abstract
We focus here on Deep Research Agents (DRAs), which are systems that can take a natural language prompt from a user, and then autonomously search for, and utilize, internet-based content to address the prompt. Recent DRAs have demonstrated impressive capabilities on public benchmarks however, recent research largely involves proprietary closed-source systems. At the time of this work, we only found one open-source DRA, termed Open Deep Research (ODR). In this work we adapt the challenging recent BrowseComp benchmark to compare ODR to existing proprietary systems. We propose BrowseComp-Small (BC-Small), comprising a subset of BrowseComp, as a more computationally-tractable DRA benchmark for academic labs. We benchmark ODR and two other proprietary systems on BC-Small: one system from Anthropic and one system from Google. We find that all three systems achieve 0% accuracy on the test set of 60 questions. We introduce three strategic improvements to ODR, resulting in the ODR+ model, which achieves a state-of-the-art 10% success rate on BC-Small among both closed-source and open-source systems. We report ablation studies indicating that all three of our improvements contributed to the success of ODR+.
1  Introduction
In this work we focus on the problem of developing Deep Research Agents (DRAs), wherein our goal is to develop a system that can take as input a natural language prompt from a user, then autonomously search for and utilize internet-based content to address the prompt. This is a challenging problem because, in principle, it typically comprises several sub-problems that are each challenging for contemporary artificial intelligence methods: for example, breaking a natural language prompt into easier sub-questions, reasoning about the use of an internet search engine to find relevant information on the internet, and then reasoning about that retrieved content to address the original prompt. Until recently there has been relatively limited progress in this area due to the difficulty of the problem, however large language models (LLMs) have demonstrated the potential to address many of these challenges and several organizations have developed proprietary systems that seek to perform Deep Research. Examples include OpenAI’s recent Deep Research [3] and Google’s Deep Research [6], Perplexity’s research capabilities [2].
One challenge with LLM-based DRAs is evaluating their performance. Benchmark problems for DRAs should simultaneously satisfy two competing properties. First, the problems must ideally be sufficiently challenging so that they cannot be easily addressed by existing methods, such as a single prompt to a LLM, or a simple single query to a browser (made either by a human or an LLM). Given the impressive capabilities of LLMs in recent years, the difficulty of problems that are suitable for DRAs has increased rapidly. Some existing benchmarks are potentially suitable for DRAs, such as HotpotQA [5] and Natural Questions [17], however recent LLM-based methods have achieved near-perfect accuracy, motivating the need for more challenging benchmark problems. Building such a benchmark is challenging however, because any benchmark questions should also include ground truth solutions, and also importantly, it should be possible to find these solutions on the internet. Without this latter property, the benchmark problems cannot be solved by a DRA. Therefore DRA benchmark problems must simultaneously be so difficult that their solutions are difficult to find, but we must also be certain that there is a solution that can be found on the internet.
Very recently, the BrowseComp [4] benchmark was introduced to address the limitations of existing benchmarks. BrowseComp includes over 1200 problems that are, by design, challenging to solve both for humans and existing LLM-based systems, while also being highly likely to have solutions that can be found via internet search. The authors of BrowseComp [4] benchmarked several proprietary systems, finding that most existing DRAs performed poorly, achieving less than 10% accuracy. The very best system, which utilized specialized methods, and substantial test-time compute, was able to achieve 50% accuracy. One major limitation of BrowseComp however is that it focused entirely on proprietary closed-source DRAs from OpenAI, creating a limited understanding of existing DRA capabilities, and a significant barrier for researchers seeking to understand, evaluate, and advance DRAs. At the time of this work, there was only one open-source DRA, termed Open Deep Research (ODR) [8]. However, the performance of ODR has yet to be quantitatively evaluated, making it unclear how open source DRAs compare with proprietary counterparts, and there is currently no baseline methods upon which to improve DRAs within the open research community.
Contributions of this work. To address these limitations, we develop BrowseComp-Small, a more computationally tractable deep research benchmark, comprising two disjoint sets of sixty questions sampled from BrowseComp: a training set, intended for DRA development; and a testing set, intended for DRA performance evaluation. We found that ODR is unable to answer any of the challenging questions in the BrowseComp-Small testing set. We propose several methodological improvements to the ODR system to support more effective deep research, resulting in our proposed ODR+ system. ODR+ successfully answers 20% of the training and 10% of the test BrowseComp questions, and therefore greatly outperforms the original ODR system. We also find that ODR+ outperforms several proprietary closed-source systems as well. Using ablation studies on the BrowseComp benchmark, we demonstrate the benefit of each of our proposed methodological improvements to ODR. We provide an open-source implementation of ODR+ [^1] https://github.com/EngDoLabadi/odrplus. We summarize our contributions as follows:
  1. 1.
    We present one of the first quantitative benchmarks of open (or closed) DRAs, and the first such benchmark on the challenging recent BrowseComp benchmark.
  2. 2.
    We introduce ODR+, an open-source DRA that achieves state-of-the-art performance on the BrowseComp benchmark among open-source DRAs. We release the code for ODR+ to support continued progress.
  3. 3.
    We present ablation studies that provide evidence of the effectiveness of our individual proposed methodological innovations, providing insights to the community on building more effective DRAs.
We wish to note that there is rapid progress on DRAs, and several open DRA benchmarks were published very recently, during the course of our work; we discuss these in Sec 2. The remainder of this paper is organized as follows: Sec. 2 discusses related work; Sec. 3 discusses the BrowseComp benchmark; Sec. 4 discusses ODR+, including a review of the original ODR system; Sec. 5 details our methodology; Sec. 6 discusses our experimental results; and Sec. 7 discusses our conclusions.
2  Related Work
Deep Research Agents. Early work in autonomous deep research began with WebGPT (2021) [1], which allowed LLMs to issue real Bing search queries and cite results. However, it supported only single-turn QA and was not released as a reusable browsing agent, making it unsuitable for benchmarking here. Recent proprietary systems represent a major advance in apparent capability, which has recently been substantiated through some empirical benchmarking (discussed below). Notable examples include OpenAI’s Deep Research, Google Gemini Deep Research, and Perplexity AI, which demonstrate varying degrees of real-time search and multi-step information retrieval. Despite their impressive performance however, these systems are closed-source, motivating the need for open, analyzable alternatives.
Existing Open Deep Research Agents. A growing number of open-source systems have recently emerged with the goal of replicating the capabilities of proprietary deep research agents. To our knowledge, the earliest and most relevant example is Open Deep Research (ODR), which was limited due to its lack of benchmarking, and therefore motivated our work here. During preparation of our work, very recently, several other open deep research agents have been published.
DeepResearcher [14] introduced a reinforcement learning framework for training browsing agents that autonomously decide what to search, read, and extract from the web. Another recent system is WebThinker [13], proposed a modular architecture to interleave deep Web exploration with reasoning, focusing on scientific and factual question answering. Although both provide code bases, their implementations are complex and are not easy to reuse or modify. Due to our limited computational budget and their very recent introduction, we did not include these methods in our benchmarking.
Existing Deep Research Benchmarks. Multi-hop reasoning across sources is a key challenge for Deep Research agents. Benchmarks such as HotpotQA [5] and 2WikiMultiHopQA [9] focus on 2-hop questions but operate within a closed setting, where the agent must find answers using a fixed set of Wikipedia passages. These benchmarks do not require open-ended web search or the ability to generate and revise search queries. As a result, they do not test whether systems can autonomously generate effective queries, plan multi-step retrieval strategies, or identify when additional information is needed, capabilities that are central to Deep Research agents. By contrast, the recent BrowseComp benchmark [4] is explicitly designed to evaluate complex, multi-hop question answering that requires synthesizing information from multiple web sources. BrowseComp questions require issuing multiple web searches, retrieving diverse pieces of evidence, and synthesizing information across documents, making it a more appropriate benchmark for studying Deep Research agents.
In addition to BrowseComp, other recent benchmarks have further advanced the evaluation and design of Deep Research agents. Mind2Web 2 [11] introduces 130 long-horizon tasks that require agents to browse unfamiliar websites and generate structured, cited answers. It also proposes an ”agent-as-a-judge” framework where agents evaluate the factuality of their own outputs. However, the benchmark emphasizes self-assessment rather than complex information synthesis and was released too recently for inclusion in our study.
Deep Research Bench [12] focuses on 89 live web tasks that test retrieval and synthesis across open-ended queries. It provides reference answers and clear evaluation criteria to assess information accuracy and completeness, but it covers fewer domains and question types than BrowseComp, with an emphasis on factual verification and summarization rather than multi-hop or exploratory reasoning. Huang et al. [10] provide a broad taxonomy and analysis of Deep Research systems, offering a high-level overview of how recent benchmarks and agent designs relate to one another.
3  The BrowseComp-Small Benchmark
We first introduce the BrowseComp (BC) benchmark, from which we create the BrowseComp-Small (BC-Small) benchmark that we use in our experiments to evaluate competing Deep Research systems. BrowseComp. [4] is a benchmark for Deep Research agents comprising 1,266 questions covering a wide range of topics, including entertainment, science, history, politics, and geography. BrowseComp questions were constructed so that they are challenging to solve, but to have short answers that are easy to verify. Questions were constructed by human ”trainers” using the following procedure. Each question was constructed by first identifying some object (e.g., a person, place, or thing), and then selecting a set of properties about that object that would, collectively, uniquely identify it. Using the selected set of properties, the designer would construct a question that asks the DRA to identify the object that satisfies all of the selected properties. For example, one BrowseComp question asks: ”Which 90s TV series starred an actor born in Tennessee, an actor who was a Caribbean immigrant, and an actor whose father was a law enforcement officer for more than 3 decades? The series was short-lived.”. The authors of BrowseComp used various criteria to ensure the difficulty of each question. For example, human trainers ensured that each question could not be correctly answered by another person within ten minutes. They also confirmed that existing models such as ChatGPT (with and without browsing) and an early version of OpenAI’s Deep Research were unable to solve them. See [4] for full design details.
BrowseComp-Small (BC-Small) The computational cost of evaluating DRAs on the full BrowseComp benchmark is large, especially for academic labs. To make BrowseComp more accessible while maintaining its utility as a benchmark, we created a smaller benchmark - termed BrowseComp-Small (BC-Small) - that comprises a subset of 120 questions from BrowseComp. We selected the questions to maintain a similar distribution of topics. Crucially, and in contrast to BrowseComp, we split BC-Small into 60 questions that are used for DRA development - essentially a training set - and another disjoint 60 questions as a validation set. This should help ensure generalization. Our choice of 120 questions was chosen to be consistent with the size of other recent public benchmarks, such as Deep Research Bench [12] (89 questions) and Mind2Web2 [11] (130 questions), where the authors also cited the high cost of issuing multiple search queries, parsing content, and performing iterative reasoning for each example.
4  Open Deep Research
Here we describe the ODR system from [8], upon which our proposed ODR+ is based. Here we provide the essential system-level details of its operation. An open-source implementation is also available [^2] ODR Implementation: https://github.com/nickscamara/open-deep-research..
The operation of ODR is illustrated in Fig. 1 and consists of three main steps.
(1) Web URL Search. The user submits a natural language question, which is passed to a large language model (LLM) along with a system prompt instructing it to generate a concise search query suitable for an internet search engine. The LLM produces a single general-purpose query without decomposing the question or performing intermediate reasoning.
(2) Content Extraction. ODR submits the generated query to a search engine and retrieves a list of candidate web pages. It opens the top-ranked link and extracts the rendered (i.e., human-visible) content of the page, which is converted to plain text and passed forward without additional filtering or parsing.
(3) Response Generation. The extracted text and original user question are then passed to the LLM, along with a prompt instructing it to generate an answer using only the retrieved content. The model produces a free-form response—a natural language answer that is not constrained to a fixed structure—which is returned to the user.
Figure 1: ODR system architecture illustrating the three main steps: (1) Web URL Search; (2) Content Extraction; and (3) Response Generation.
Figure 2: ODR+ system architecture, illustrating three major steps: (1) Question decomposition; (2) an iterative Sub-solution Search step, which seeks internet-based evidence to address each sub-question; and (3) Response Synthesis, where a structured response, denoted $r_{+}$, is generated for the user based upon a summary of evidence from the internet.
5  Open Deep Research Plus (ODR+)
Here we describe our system, ODR+, which is constructed by making several improvements to the ODR system. ODR provides an important initial working system, but it suffers from several limitations that cause it to fail on complex, multi-hop research questions. We hypothesize that ODR fails on these problems for at least three reasons: it does not decompose the user query into simpler sub-questions; it lacks any form of iterative reasoning or adaptive planning; and it is not prompted to produce structured output. ODR+ addresses these limitations through the introduction of three modules illustrated in Fig 2: Question Decomposition, Sub-solution Search, and Response Synthesis. We next describe each of these three major modules, as well as sub-modules that contribute to them, which are detailed in pseudocode for ODR+ in Algorithm 1. The engineered prompts used throughout the ODR+ pseudocode are summarized in Table 1.
Prompt Prompt Name Description
$P_{1}$ Constraint Extraction Ask the LLM to extract the most specific constraints (e.g., names, dates, descriptors) from the user query to guide downstream reasoning.
$P_{2}$ Sub-question Generation Instruct the LLM to reformulate the original query into focused sub-questions that preserve key constraints.
$P_{3}$ Content Extraction Direct the LLM to extract only facts from retrieved web content that match specified constraints and current sub-question.
$P_{4}$ Evidence Analysis Ask the LLM to evaluate current findings, determine sub-question completion, and propose new sub-questions or termination.
$P_{5}$ Response Synthesis Instruct the LLM to aggregate all findings and output a structured final answer with confidence and justification.
Table 1: Summary of engineered prompts used in ODR+ system, ordered by execution sequence.
Question Decomposition
The first module of ODR+ converts the original user query, $userQuery$, into a set of focused sub-questions, as detailed in lines 7-8 of Algorithm 1. This begins with a call to extractConstraints (Line 7), which takes as input the prompt $P_{1}$ in Table 1 and the original query. The prompt instructs the language model to extract highly specific identifying constraints—such as names, dates, locations, and numerical values—that help narrow the answer space. The output is a structured JSON array of constraints.
Next, the system calls generateSubQuestions (Line 8), which receives $P_{2}$, the original query, and the extracted constraints. This prompt guides the model to generate a small number of clear, fact-based sub-questions that target the extracted constraints. The resulting sub-questions are stored in the queue S.subquestions, which forms part of the system’s internal research state S. This state also tracks retrieved evidence, depth of search, processed URLs, and intermediate results, as initialized in Lines 4–5.
Iterative Sub-Solution Search
This module focuses on addressing each of the sub-question identified in module (1) and is shown in lines 10–30 of Algorithm 1. The iterative process continues until all sub-questions are addressed, or some other stopping criteria is met (e.g., permissible run-time, denoted $T_{max}$, is exceeded; or a maximum number of sub-questions, denoted $D_{max}$ is exceeded).
At the beginning of each iteration, the system selects an unresolved sub-question from S.subquestions (Line 12) and uses it directly as a web search query. We observed that web browsers returned a different page ranking each time the same query was submitted and therefore, to reduce variability, we submitted the same query $N_{\text{query}}$ times using webSearch (Line 13). From each of the $N_{\text{query}}$ search attempts, the top-ranked URLs are collected. The frequency of each URL across all attempts is then tallied, and the $k$ most frequently occurring URLs are selected using selectMostFrequent (Line 14) for further processing.
Next, an extraction prompt extractionPrompt is constructed by invoking createExtractionPrompt with the prompt template $P_{3}$, the original user query, and the previously extracted constraints (Line 15). It is designed to instruct the LLM to extract only the portions of the page content that are relevant to the sub-question and consistent with the specified constraints. The prompt is passed, along with the full text of each selected URL, to the LLM. The model is invoked once per URL and typically returns one or more short spans of relevant text. These outputs are stored as structured findings—each consisting of the extracted text and its source URL—and are appended to S.findings, the list of accumulated findings maintained in the internal research state (Line 17).
After collecting new findings, ODR+ invokes the language model using analyzeEvidence with prompt $P_{4}$, the current sub-question, and the full set of accumulated findings (Line 18). The prompt instructs the model to produce a structured response in JSON format, including fields such as a proposed answer to the sub-question (if one can be determined), a confidence score, a list of satisfied constraints, and any suggested follow-up sub-questions. The JSON output is parsed, and if a valid answer is provided, it is added to S.subAnswers (Lines 19–20). If follow-up sub-questions are suggested, they are appended to S.subquestions (Lines 21–22). Additionally, the control flow logic uses the model’s analysis—specifically the confidence score and recommendation on whether to continue—to decide whether to proceed to the next iteration or terminate the loop early (Lines 23–24).
Algorithm 1 Open Deep Research Plus (ODR+)
1:Input: userQueryuserQuery (original user question)
2:Output: structuredResponsestructuredResponse (formatted final answer)
3: {Module 1: Question Decomposition}
4:Initialize:
[…]
9: {Module 2: Iterative Sub-Solution Search}
10:while $S.depth<D_{max}$ and $t<T_{max}$ and     $S.subquestions\neq\emptyset$ do
11:  $S.depth\leftarrow S.depth+1$
12:  $currentSubQuestion\leftarrow S.subquestions.\texttt{pop}()$
13:  $urls\leftarrow\texttt{webSearch}(currentSubQuestion,N_{query})$
14:  $topUrls\leftarrow\texttt{selectMostFrequent}(urls,k)$
15:  $extractionPrompt\leftarrow$   $\texttt{createExtractionPrompt}(P_{3},userQuery,constraints)$
16:  $newFindings\leftarrow\texttt{extractFromUrls}(topUrls,extractionPrompt)$
17:  $S.findings\leftarrow S.findings\cup newFindings$
18:  $analysis\leftarrow$   $\texttt{analyzeEvidence}(P_{4},S.findings,currentSubQuestion)$
19:  if $analysis.subAnswer\neq null$ then
20:   $S.subAnswers\leftarrow S.subAnswers\cup\{analysis.subAnswer\}$
21:  end if
22:  if $analysis.subquestions\neq\emptyset$ then
23:   $S.subquestions\leftarrow$   $S.subquestions\cup analysis.subquestions$
24:  end if
25:  if ($analysis.hasAnswer\land analysis.confidence\neq low$) or     $\neg analysis.shouldContinue$ then
26:   break
27:  end if
28:  wait($W_{ms}$)
29:  $t\leftarrow\texttt{getCurrentTime}()-startTime$
30:end while
31: {Module 3: Response Synthesis}
[…]
33:return  structuredResponsestructuredResponse
By iteratively resolving sub-questions in a goal-directed manner and updating its internal state at each step, ODR+ constructs a coherent body of evidence grounded in retrieved content, enabling robust and explainable research over complex factual queries.
Response Synthesis
The third and final module of the ODR+ system—implemented in lines 32–33 of Algorithm 1—is responsible for synthesizing the final structured answer once the iterative search loop terminates.
On line 32, the system invokes the language model using the engineered prompt $P_{5}$, along with the original user question (userQuery), the extracted constraints (constraints), and the accumulated evidence (S.findings). The prompt instructs the model to generate a final answer grounded strictly in this structured content, explicitly forbidding reliance on prior knowledge. The model is asked to produce a response in the standardized BrowseComp format:
Explanation: {reasoning based on findings}
Exact Answer: {short final answer or ’Unknown’}
Confidence: {confidence score as a percentage}
The prompt also instructs the model to compute a confidence score based on the number of key constraints satisfied by the proposed answer, relative to the total number of extracted constraints.
On line 33, the model’s output is stored in structuredResponse. The system then validates this response to ensure that all required fields are present and correctly formatted. If any field (such as the explanation, exact answer, or confidence score) is missing or malformed, fallback values are inserted. For example, the system may assign "Unknown" as the answer and a default confidence score of 10%.
This validation step ensures that every final output is complete, properly structured, and ready for automated evaluation. By grounding all answers in retrieved evidence and enforcing a consistent output format, ODR+ produces responses that are reliable, transparent, and evaluation-ready.
6  Numerical Experiments
We conduct experiments on our BrowseComp-Small benchmark (see Sec. 3), which comprises two disjoint sets of sixty questions: a training set, and a testing set. We evaluate several competing DRA systems on the sixty test questions: ODR, ODR+, Claude-DeepResearch (Anthropic) [15], and Gemini-DeepResearch (Google 2025) [16]. We used the sixty questions in the BrowseComp-Small training set to develop ODR+.
ODR+ Development and Hyperparameter Settings
All development of ODR+ was done using the sixty training questions in our BrowseComp-Small benchmark. This was done to minimize the potential of overfitting the design of ODR+ to the testing questions. Many steps of ODR+ (and ODR) utilize an LLM, and we utilized the GPT-4o-mini model via the OpenAI API. This model was chosen for its favorable balance of cost, latency, and reasoning quality, enabling scalable evaluation under limited compute budgets. We utilized the following hyperparameter settings for ODR+:
  • Search Depth ($D_{\text{max}}=6$): The system performs up to six iterative search hops per question.
  • Time Limit ($T_{\text{max}}=210$ seconds): Each question must complete within 3.5 minutes of wall-clock time.
  • Top-$k$ URLs ($k=3$): At each hop, the system selects the $k$ most frequent URLs across multiple search attempts.
  • Search Retries ($N_{\text{query}}=3$): Each sub-question is submitted to the search engine $N_{\text{query}}$ times to reduce variability in returned results.
These hyperparameters were chosen through experimentation on the training set, balancing answer quality, runtime, and the cost of running the model. We note however that increasing these hyperparameter settings may likely improve system accuracy, at the cost of increased computational cost — we did not have the resources to investigate this potentiality.
Evaluation Methodology
We follow the official BrowseComp evaluation protocol, which requires system responses to conform to a standardized three-part structure (Explanation, Exact Answer, Confidence). Each system output is scored using the released BrowseComp evaluator, which leverages the GPT-4o model (via OpenAI API) to assess both answer correctness and formatting adherence. The evaluator performs semantic comparison between the predicted answer and the ground truth to determine exact match accuracy. Therefore, for each question the evaluator determines whether the response of the DRA is correct or incorrect, and we report the resulting accuracy over the 60 test questions of each system. All web searches and page extractions in ODR and ODR+ were performed using FireCrawl [7] to ensure consistent and structured retrieval.
Main Results
The main results are reported in Table 2, where the results indicate that the ODR system was unable to answer any of the questions in the BrowseComp-Small test set. By contrast, ODR+ answered 10% of test questions (6 of 60) with exact-match correctness. To our knowledge therefore, ODR+ achieves the current state-of-the-art (SOTA) performance on the BrowseComp benchmark, among open-source models. Surprisingly however, we found that ODR+ also outperformed the other two proprietary DRAs that we evaluated on BrowseComp-Small: Claude-DR and Gemini-DR. At the time of our testing, Claude-DR and Gemini-DR both returned 0% exact match accuracy on our test set. In our manual evaluations, we observed that these systems often produced long reports rather than concise answers - an approach that is penalized under BrowseComp’s strict evaluation criteria. By contrast, ODR+ had the advantage that we designed it specifically to solve BrowseComp questions, and therefore it is designed to focus on satisfying the exact constraints of the questions, and generating structured verifiable answers.
For completeness, we also report the results of ChatGPT-DR that were published in [4], which were obtained on the full BrowseComp benchmark. We were unable to evaluate ChatGPT-DR ourselves because it lacked a public API. In [4] the authors reported different performance depending upon the amount of test-time compute that was applied, ranging from approximately 10% with limited test-time compute, up to 51.5% with extensive test-time compute. The precise amount of test-time compute is not provided, making it difficult to judge precisely how this system would compare with ODR+, given the same computational resources. Notably, in contrast to all other systems we consider, the ChatGPT-DR system reported in [4] was developed using the entire BrowseComp benchmark, and without utilizing a disjoint training and testing set, potentially providing it with an advantage.
Table 2: Performance Comparison on BC-Small Test Set
Deep Research Agent LLM Accuracy (%)
ODR GPT-4o-mini 0%
ODR+ (ours) GPT-4o-mini 10%
Claude-DR Sonnet 4 0%
Gemini-DR Gemini 2.5 Pro 0%
ChatGPT-DR GPT-4o $\sim$10–51.5%*
*Results reported from [4] on full BrowseComp Benchmark
Additional Analysis and Ablations for ODR+
In this section we report some additional analysis of ODR+. We first discuss some additional performance metrics for ODR+ that provide insight on its functionality. First we report that ODR+ obtained 20% accuracy on the training set, which is twice as high as we obtained on the testing set. The size of our testing and training set are relatively small, but this could also indicate some degree of overfitting of our system to the training set; this remains an open question. We also found that 25% of test questions (15 of the 60) exceeded the time limit ($T_{max}$), and were terminated before completion, highlighting the potential positive impact of increasing test-time computation. Despite this, 98.3% of all test questions (59/60) produced structurally valid outputs—containing a well-formed Explanation, Exact Answer, and Confidence field—demonstrating strong robustness across system components even in partially completed runs. The average runtime across the 45 successfully completed test questions was 198 seconds (out of a 210-second limit), with longer runtimes typically observed in questions involving more search hops or fallback retries.
To understand the impact of individual components in ODR+, we conducted ablation studies by disabling key modules and observing performance changes. Due to computational costs, we randomly selected 20 test questions from the BC-Small test benchmark and evaluated the following ablated variants of ODR+:
  • No Sub-question Decomposition: This variant disables the sub-question generation and decomposition step. The system reverts to a single-query approach similar to the original ODR. We expect a significant drop in performance on multi-hop questions, which typically require breaking down complex prompts into smaller, focused searches.
  • No Iterative Planning: This variant removes adaptive planning and research state management. Sub-questions are processed sequentially, without retry logic or feedback based on earlier findings. This limits the system’s ability to adjust its strategy dynamically, likely reducing the effectiveness of information gathering.
  • No Structured Synthesis: This variant disables structured output formatting and validation. Instead, it uses free-text generation similar to the original ODR. We anticipate lower confidence estimates, formatting inconsistencies, and a higher likelihood of inaccurate or incomplete final answers.
Table 3 shows how disabling each core module reduces the accuracy of the ODR+ system, highlighting the overall contribution of each component to system effectiveness.
Table 3: Ablation Study Results on 20-Question Subset
System Variant Accuracy (%)
ODR+ 25% (5/20)
No Structured Synthesis 0% (0/20)
No Sub-question Decomposition 5% (1/20)
No Iterative Planning 5% (1/20)
7  Conclusions
We introduced ODR+, an enhanced open-source Deep Research Agent (DRA) designed to perform complex multi-hop web-based question answering. Building on the original ODR system, ODR+ incorporates sub-question decomposition, iterative planning, and structured synthesis, enabling more effective and interpretable autonomous research across open-domain queries.
We benchmarked ODR+ on the BrowseComp-Small dataset and demonstrated that it significantly outperforms the original ODR baseline, achieving 10% exact-match accuracy on the test set while maintaining high structural validity. Our ablation studies confirmed the critical role of each added component in improving performance.
To support continued progress in the development and evaluation of DRAs, we release our implementation and tools publicly. We hope ODR+ serves as a foundation for future research in open, analyzable, and extensible Deep Research Agents.
Acknowledgements
The authors would like to thank WattTime and the Climate TRACE coalition for their organizational support, and Climate TRACE funders – Al Gore, Benificus Foundation, Clean Air Fund, Google.org, the partners of Generation Investment Management, Holdfast Collective, the Patrick J. McGovern Foundation, and Schmidt Futures.